OpenStudy (anonymous):

Length = 5x_4 width = x+3 height =2x_1 Find volume using V=lwh

OpenStudy (anonymous):

vol = (5x-4)(x+3)(2x-1) = [5x^2 +15x -4x - 12](2x-1) = [5x^2 + 11x -12](2x-1) = 10x^3 - 5x^2 + 22x^2 - 11x - 24x + 12 = 10x^3 + 17x^2 - 35x + 12

OpenStudy (anonymous):

ur good!